EVA Report – May 9th

EVA #04

Author: Camilo Andrés Zorro Mendoza

Purpose of EVA: Second testing of drones and first drones flight and thermal rover tests.

Start time: 08:00 AM

End time: 09:45 AM

Narrative:

Today we were running two experiments, following the debriefing proposed yesterday we implemented the recommendations of the buddy system. After pressurization process, a walk was done around the Hab in order to select some points of interest for observe it in a safe distance. After, it was making the pressurization process of the RAM and the drones were taking.

Two experiments were running at the same time, in the first experiment a thermal blanket was tested in order to find the most appropriate number of layers to decrease the heat transfer radiation effects without increasing the effects from conduction, for that was made 5 different configurations. The second one two crewmembers tested the takeoff and landing of a fixed wind vtol drone and multicopter drone. A number of lessons were learnt and are summarized below from the after EVA debrief.

– All radios should be at the channel 12 for the EVA operations, communication with the Hab is critical.

– It was clarified that the radios in the Hab should be in the channel 9. Channel 12 it is only when it is a EVA operation in process.

Destination: Hab surroundings

Coordinates (UTM NAD27 CONUS): Sector12S Northern 4250933, Eastern 518167

EVA Participants: Zoe, Vittorio (Medic), Paolo and Camilo (Commander)

Road(s) and routes per MDRS Map: None

Mode of travel: walking

Operations Report – May 9th

SOL: 4

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night):

On (last night): 8:00 PM

Off (this morning): 7:00 AM

Solar— SOC 85% at 7:20 PM

Diesel Reading – 97%

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 230 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 160 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143406.8

Toilet tank emptied: No

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 92.1h

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours: 58.7h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 94.3h

Beginning charge:

Ending charge:

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: Nothing to report

Summary of internet: Signal was down during the morning, then returned after lunch.

Summary of suits and radios: Some repairs were done on the suits.

Summary of Hab operations: A fantastic pesto sauce was prepared with the basil from the GreenHab

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: The damaged autoclave has been successfully used for heating (the pressure functions are unavailable). [From Director]: the binocular scope is damaged. I used it this evening and it will need some adjustment and

hopefully I will be able to fix it. I will try tomorrow morning.

Summary of RAM operations: It was depressurized to allow the EVA crew to take the drones for the experiments.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: The water heater is on, but it seems to be not functioning properly. Is there a procedure we are not doing correctly?

Science Report – May 9th

Crew 212 – LATAM III

Submitted by GreenHab Officer Hector Palomeque

1. Solar Observation/Astrophotography/Stellar Clusters:

• The eyepiece of MUSK Telescope was successfully assembled.

• First picture of the sun with MUSK Solar Telescope (a bit cloudy).

• Deep Sky Robotic Telescope still observing a star cluster (interrupted due to the bad weather).

2. X-5:

• The first flight of 2 on 3 drones in a Martian environment was performed during today’s EVA (SOL 4).

• Manual light off and landing did for checking that Drones are working correctly.

• Drones operation with space suit was tested.

• Waiting for the weather forecast to schedule and perform the next experiment (photos of the MDRS Hab and other areas).

3. VESTA:

• One flight of the Drone-Multicopter were performed around MDRS Hab during today’s EVA (SOL 4)

• Interesting points around MDRS Hab selected.

• Drone successfully connected to the computer for planning an automatic mission.

4. Observation Rover:

• Linux system correctly launched and running in a crew’s Laptop.

• Still needing to download the script from UK for uploading it into the Raspberry Pi Rover board.

5. Space Farming:

• Red-Clay collected for preparing a mix for Martian-like regolith.

• Broken-Autoclave used for boiling in order to disinfect/sterilize the material as much as possible (without problem nor danger).

• Broth culture media for selected Chernobyl and Fukushima bacteria successfully prepared and sterilized (no contamination detected in negative control after 24 hours).

• Chernobyl and Fukushima bacteria correctly inoculated in Nutrient Broth.

• Chernobyl and Fukushima bacteria already showing sights of growth.

6. Keep Talking:

• Crewmembers getting a better performance during the game. Some teams already have reached the Difficult Level.

• The fourth day of playing (corresponding for each SOL).

7. Ethnography study:

• Audio-recordings with the information collected during the observations of the Crew 212 behave already sent to the Colombian Aerospace Research and Develop Group (Colombia).

8. Pulse Oximeter Measurement:

• No update to report.

9. Software for macroscopic characterization:

• Nutrient Agar in Petri dishes prepared using the Broken-Autoclave (all correct).

• MDRS soil bacteria cultivated in Nutrient Agar.

• Waiting for bacterial growth (24 hours).

• EVA scheduled for collecting rhizospheric soil of MDRS plants.

Science Report – May 9th

Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 4

Sim suit used: 3, 4, 8 and 9.
used for 2 hours

Sim suit: 3
Voltage level before EVA: 13.4 V
Voltage level after EVA:
12.8 V

Sim suit: 5
Voltage level before EVA: 13 V
Voltage level after EVA:
12.7 V

Sim suit: 8
Voltage level before EVA: 13.2 V
Voltage level after EVA:
10.5 V

Sim suit: 4
Voltage level before EVA: 12.9 V
Voltage level after EVA:
10.6 V

Journalist Report – May 9th

Today we woke up to start with a morning EVA. Two experiments were run in tandem, this was the rover and drone project. The rover project was the retesting of the thermal blanket, with assistance, as per the debrief of EVA#4. The initial results suggest a successful experiment. In addition two of the three drones went for their first flight test. This involved first finding points of interest around the Hab, these points will be explored in future EVA with drones.

First observations using the solar telescope were undertaken by two crew members looking through the eye piece. Next time there is hope that the pictures can be taken of the sun.

Secondly, the problem solving experiment was ran twice. This was to make up for the missed day on Sol 1 where the crew misunderstood the explanation of the game.

After a successful European Cultural night where pasta, cheeses and desserts of Europe were showcased the left over pasta was eaten today. Basil was harvested from the GreenHab and used to make homemade pesto for the pasta. The carrots will be used tonight for the Latin American Cultural night as part of the Peruvian dishes. We love the Greenhab!

Interviews were compiled for the WOMAD festival science tent. These videos will be used as part of a final video presentation that will be displayed at the festival. In addition, written statements were taken for The Engineer blog.

GreenHab Report – May 9th

GreenHab Officer: Hector Palomeque

Environmental control: Ambient with the door open from 07:00 to 15:00 + door closed from 15:00 to the rest of the day

Shade cloth 80%: on

Average temperatures:

Low: 15.6 °C (±1.19)

High: 35.0 °C (±1.28)

Hours of supplemental light: 0

Daily water usage for crops: 2 gallons + 2 gallons (+ 3 gallons for ScienceDome)

Water in Blue Tank: 162 gallons

Time(s) of watering for crops: 07:00 + 18:00

Changes to crops: All plants look healthy. Chard is ready to harvest. Basil is still blooming. Tomatoes in the black pots start to bloom. The area that was occupied by carrots now is empty for any other crop (Recommendation: herbs, tomato or any other with aerial growth/fructification. Tubers tends grows abnormally small due to the lack of space for dig).

Narrative: The GreenHab door had to been closed at 15:00 due to the rain and wind. Plants were watered two times with half of the usual volume because the substrate remains moisture longer thanks to the cold weather.

Harvested: 5g Sage + 59g Basil

Support/supplies needed: None.

Astronomy Report – May 9th

Name: Mariona Badenas

Crew: 212

Date: 9 May 2019

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested (choose one): MDRS-WF

Object Viewed #1: IC5067

Problems Encountered: The telescope took pictures of this nebula for some time until the weather became too bad.

(More objects to be Viewed): NGC6992 and the Veil Nebula

Problems Encountered: N/A (Observations of these objects have not begun yet).

MUSK OBSERVATORY

Solar Features Observed: The crew engineer and I managed to observe the sun through the eyepiece (or at least that’s what we think). We also attempted to take pictures, but it was cloudy (see the attached photograph).

Problems Encountered:

1. We eventually had to stop observing due to bad weather (clouds).

2. The telescope’s pointing system was not functioning properly, so I had to use the telescope manually. I have attached a picture of the telescope once the alignment and pointing process is complete, and you’ll find that it is not pointing at the Sun.

3. Given that the pointing was not correct, I couldn’t verify whether the tracking system worked. In the event that it didn’t, my plan of action will be to follow the procedure that Peter described to Crew 211 last week.

4. This is a question, rather than a problem: what is the time system that the telescope uses for its own measure of time (h:m:s)? I assumed it was the Standard Time in Utah, but I was wondering whether this assumption is incorrect. Inserting a wrong time in the telescope would affect its pointing.

Sol Summary – May 9th

Sol: 4

Summary Title: It’s raining martians!

Author’s name: Marlen Castillo Vilcahuaman

Mission Status: After some rain, we’re ok.

Sol Activity Summary: We had a morning EVA today. Weather was nice for experiment. Until this day, all of the crewmembers have some advancement in their experiments. We’ll be sending a Science Report today.

Look Ahead Plan: 2 EVAs for tomorrow and, hopefully, more good news for our experiments.

Anomalies in work: None.

Weather: Rainy at the afternoon.

Crew Physical Status: Still a crew member with cold sore.

EVA: 1

Reports to be file: Operations Report, GreenHab Report, Journalist Report, Daily Photo Report, EVAs Request, Space Suit Report, Astronomy report, Science Report

Support Requested: The Non-Burnable trash bag needs to be picked up. Also, the owl bag needs to be retired from the airlock.